Sodium carbonate can be made by heating sodium bicarbonate:
2NaHCO3(s) → Na2CO3(s) + CO2(g) + H2O(g)

Given that ΔH° = 128.9 kJ and ΔG° = 33.1 kJ at25°C, above what minimum temperature will the reaction becomespontaneous under standard state conditions?
Answer #1
We know that, ΔG0 = ΔH0- TΔS0
  ΔH° = 128.9 kJ
  ΔG° = 33.1 kJ
   T = (25 + 273) K = 298 K
   ∴ 33.1 kJ = 128.9 kJ - 298(ΔS0)
        ΔS0 = 0.321 kJ /K
The temperature above which the reaction is spontaneouscan be obtained by setting ΔG°= 0
             ΔG0 = ΔH0 -TΔS0
                 0 = ΔH0 - TΔS0
                  T= ΔH0 / ΔS0
                    = 128.9 kJ / 0.321 kJ /K
                    =401.6 K
At a temperature above 401.6 K, TΔS0 hasa larger magnitude than ΔH0 andtherefore,  ΔG0 becomesnegative.
Thus, the reaction becomes spontaneous at 401.6 K